# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Notes:
This table summarizes the statistics on the sales in the retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
@@
2. The number of stores reported the share of regular customers (%) are 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively.
@@
Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
In Table C.2, I analyze the relationship between credit purchases and competitiveness among local retailers using the selling item by scale of business cells reported in Table C.1. In Column (1), I regressed the sales in credit purchases (in percentage points) on the number of local competitors. I then conditioned the business scales in Column (2). The $F$-statistic $p$-value for joint zero-slope hypotheses for these scale dummies is estimated to be $0.990$. In Column (3), I included the selling item dummies instead of the scale dummies. The $F$-statistic $p$-value for joint zero-slope hypotheses is rejected at the conventional level, and the measure of fit is relatively high compared with the other two models reported in Columns (2) and (3). The estimate is $3.420$, suggesting that a one unit increase in the local competitors increases the use of credit purchase by roughly $3.4$ percentage points. Given that the standard deviation of the local competitors is approximately $2.3$, the estimated magnitude corresponding to the one standard deviation increase would be $7.9$ ($3.420\times 2.3$). This is roughly $41$% of the standard deviation of the dependent variable ($19.2$) and thus, is an economically meaningful magnitude.
@@
C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
Table C.3 summarizes the use of credit purchase and the losses on the credit sales documented in the 1930 Survey Report. Following the report’s style, the item categories are divided into six categories: white rice, other grains, vegetables, fruit, fish, and fuel. The total number of retailers in each category is $157$, $7$, $67$, $16$, $46$, and $64$, respectively. As explained in Section 5, most of them used credit purchases as their selling method. Importantly, among the retailers using credit purchases, almost all of them suffered losses in sales. For example, $99$% of white rice retailers experienced losses. The minimum value of this share is still $97$% for the fuel retailers. This means that credit purchase was a loss-making sales method for the retailers.
The statistics in Table C.3 also indicate that many cases have unknown loss amounts. Overall, however, most of the cases seem to be grouped into bins of less than $5$% of losses. For example, $87$% ($97$ retailers) of the $111$ available rice retailers suffered losses amounting to $5$% of sales. Similarly, $65$% ($32$ retailers) of the $49$ available fuel retailers are grouped into the same loss category. Systematic statistics are unavailable to indicate the percentage of total profits corresponding to these losses. However, it is worth mentioning that a large part of losses was capped by $5$% of total sales in all these different retailers. This implies that while the retailers could not avoid losses, they might have set prices that anticipated losses.
@@ C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
C.5 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
@@
Column (1) of Table C.4 presents the result for Equation 4.109109109 The estimated coefficient is close to zero and statistically insignificant. A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases. Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications. Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households that use credit for their purchases. This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.110110110Despite this, low-income households might have had more difficulty using credit for purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
@@
  • (51)
  • (52) Tanaka, Hikaru. Mouhitotsu no kinyū system (Another Financial System) [in Japanese]. Nagoya: Nagoya University Press, 2018.
@@
  • (58) Patrick, Royston and Douglas G. Altman. “Regression using Fractional Polynomials of Continuous Covariates: Parsimonious Parametric Modeling.” Journal of the Royal Statistical Society. Series C (Applied Statistics) 43 no. 3 (1994): 429–467.
@@
  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–12).
  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–13).
  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–14).
  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–15).
  • (12)

### v14

@@ C.5 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Notes: This table summarizes the statistics on the sales in the retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
@@
2. The number of stores reporting the share of regular customers (%) is 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively.
@@
Source: Tokyo Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
In Table C.4, I analyze the relationship between credit purchases and competitiveness among local retailers using the selling item by scale of business cells reported in Table C.3. In Column (1), I regressed the sales in credit purchases (in percentage points) on the number of local competitors. I then conditioned the business scales in Column (2). The $F$-statistic $p$-value for joint zero-slope hypotheses for these scale dummies is estimated to be $0.990$. In Column (3), I included the selling item dummies instead of the scale dummies. The $F$-statistic $p$-value for joint zero-slope hypotheses is rejected at the conventional level, and the measure of fit is relatively high compared with the other two models reported in Columns (2) and (3). The estimate is $3.420$, suggesting that a one unit increase in the local competitors increases the use of credit purchase by roughly $3.4$ percentage points. Given that the standard deviation of the local competitors is approximately $2.3$, the estimated magnitude corresponding to the one standard deviation increase would be $7.9$ ($3.420\times 2.3$). This is roughly 4141% of the standard deviation of the dependent variable ($19.2$) and thus, is an economically meaningful magnitude.
@@
C.6 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
Table C.5 summarizes the use of credit purchases and the associated losses documented in the 1930 Survey Report. Following the report’s classification, items are divided into six categories: white rice, other grains, vegetables, fruits, fish, and fuel. The total number of retailers in each category is $157$, 77, $67$, $16$, $46$, and $64$, respectively. These retailers used credit purchases as their primary selling method. Notably, almost all retailers using credit purchases experienced sales losses. For example, 9999% of white rice retailers reported losses, while the minimum share was 9797% for fuel retailers. This indicates that credit purchase was generally a loss-making sales method.
Table C.5 also shows that many cases have unknown loss amounts. Overall, most losses appear to be concentrated in bins of less than 55% of sales. For instance, 8787% ($97$ retailers) of the $111$ rice retailers with available data suffered losses amounting to 55% of sales. Similarly, 6565% ($32$ retailers) of the $49$ fuel retailers fell into the same loss category. Systematic statistics on the percentage of total profits corresponding to these losses are unavailable. However, a large portion of the losses is capped at 55% of total sales across all retailer types. This suggests that, although retailers could not entirely avoid losses, they may have set prices to account for anticipated losses.
@@ C.6 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
Lastly, Table C.6 summarizes the industrial and family structures in 1930 Tokyo using the official report of the Population Census. Column 1 summarizes the share of male workers in each industry, showing a similar distribution as that in 1920 (Column 1 of Panel A in Table 1). While the development of the city throughout the 1920s slightly increased the share of commercial industries, manufacturing and commercial industries covered approximately 75% (77% in 1920). Column 2 confirms that the average household size, share of married males, sex ratio, and average age of males are in a similar range to those measured in the 1920 Population Census (Column 1 of Panel B in Table 1). These show that not only the industrial structure but also the average household characteristics had not changed essentially between 1920 and 1930. While Tokyo experienced the Great Kanto Earthquake in 1923, the population structure did not change dramatically, except for a temporary evacuation soon after the earthquake.
Table C.6: Industrial and Family Structures in Tokyo: 1930 Population Census
(1) Industrial structure (2) Family Structure
Name of survey 1930 Population census 1930 Population census
Survey area Tokyo City Tokyo City
Survey subject Complete survey Complete survey
Survey month and year October 1930 October 1930
Agriculture 0.5 Average household size (in person) 5.1
Fisheries 0.1 Share of married males (%) 34.2
Mining 0.0 Sex ratio (males/females) 1.2
Manufacturing 37.2 Average age of males 25.2
Commerce 37.3
Transport 7.2
Public service and professions 13.5
Housework 0.6
Other industry 3.6
Notes: Panel A: This panel summarizes industrial structures based on the occupations of male workers recorded in the 1930 Population Census. Each share is calculated as the number of males working as regular employees in each sector divided by the total number of male workers (percentage). Unemployed males (mushoku) are excluded from these figures. Panel B: The average household size is calculated as the total number of people divided by the number of households. The share of married males is calculated as the number of married males living with/without spouses divided by the total number of males (percentage). The average age of males is calculated from the population tables by age group reported in the census. For the open-ended age category over 65, the value is rounded to $67$, following the range of the previous category (i.e., 60–64). All figures in this panel exclude a small number of quasi-households (jyun setai) such as students in dormitories and patients in hospitals. Sources: Column 1: Statistics Bureau of the Cabinet (1933, pp. 52–53). Column 2: Statistics Bureau of the Cabinet (1933, pp. 14; 6–7; 357).
C.7 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
@@
Column (1) of Table C.7 presents the result for Equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases. Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications. Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households that use credit for their purchases. This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases. [^94] Despite this, low-income workers might have had more difficulty using credit for purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
@@
  • (4) de Chaisemartin, Clement., and Xavier D’Haultfœuille. “Two-way fixed effects and difference-in-differences with heterogeneous treatment effects: a survey.” The Econometrics Journal 26, issue. 3 (2023): C1–C30.
  • (5)
@@
  • (9)
  • (10) Horrell, Sara., and Deborah Oxley. Work and prudence: Household responses to income variation in nineteenth-century Britain. European Review of Economic History 4, no.1 (2000): 27–57.
@@
  • (60) Patrick, Royston and Douglas G. Altman. “Regression using factional polynomials of continuous covariates: parsimonious parametric modeling.” Journal of the Royal Statistical Society. Series C (Applied Statistics) 43 no. 3 (1994): 429–467.
@@
  • (26) Ministry of Land, Infrastructure, Transport and Tourism (database). Gyōsei kuiki data (Administrative Area Data). [in Japanese] (https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html, accessed on August 15, 2023).
  • (27)
@@
  • (41)
  • (42) Statistics Bureau of the Cabinet. Shōwa gonen kokuseichōsahōkoku, fuken hen, dai yonkan (Population Census of Japan, Tokyo Prefecture, 1930 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1933.